# Body-Hand Modality Expertized Networks with Cross-attention for Fine-grained Skeleton Action Recognition

### v2

@@ I INTRODUCTION
Human action recognition (HAR) has garnered significant attention in robotics[1, 2], especially in areas of human-robot interaction and collaboration. The goal of HAR is to automatically identify human actions from video data—a challenging task due to various factors such as complex temporal dependencies, intricate spatial relationships, background noise, and varying camera perspectives. To address these challenges, existing studies have explored diverse data modalities such as RGB video, depth, infrared, audio, and skeleton joints.
Skeleton-based approaches, particularly those using Graph Convolutional Networks (GCNs)[3, 4] have shown great promise in efficiently modeling spatial-temporal dynamics by representing skeletal joints as graph nodes and their connections as edges. Recent architectures[5, 6, 7, 8, 9] have demonstrated strong performance in action recognition; however, most of these methods primarily focus on full-body motion analysis. This narrow focus limits their ability to capture the fine-grained hand movements essential for differentiating actions such as ”okay sign” or ”victory sign”.
@@
To address these limitations, recent research[10] extends traditional skeleton-based representations by incorporating additional hand and foot keypoints into a unified graph. To manage the increased complexity of this enriched graph, a combined pooling strategy that integrates both node pooling and instance pooling is employed in this approach. While this combined pooling method in an largely-integrated graph significantly enhances efficiency and overall recognition performance, it also results in the loss of subtle hand details, making the hand features vulnerable to blur from larger, dynamic body movements. Among the other related tasks, separating the body and hand into different graph streams is used in Sign Language Recognition[11] and Action Forecasting methods[12, 13] (see Section II for more details).
@@
Each modality is further refined in its dedicated expertized branch. Then, dividing interactive branch and expertized branch enables the two branches to exchange information without using residual connections—thereby preventing conflicts between expertized features and interactive features. Global average pooling is followed by an attention mechanism aggregating body and hand features. Integrating diversely-pooled interactive features, our approach results in a lightweight model with significantly reduced computational complexity.
@@
Inspired by MMNet [14], we leverage body skeleton features to guide the learning of RGB features. By treating body features as dynamic weights, our framework enables the RGB stream to focus on complementary contextual information, such as object interactions and environmental cues, enhancing overall recognition performance.
@@ II-A Skeleton-Based Action Recognition
Skeleton-based action recognition has evolved significantly, moving from temporal modeling approaches like RNNs, such as Long Short-Term Memory (LSTM) networks, to more advanced graph-based methods like GCNs [3, 4], CNNs[15], and Graph-based Transformers [5, 16, 17, 18, 19]. RNNs effectively modeled sequential temporal dependencies but struggled with spatial relationships. GCNs addressed these challenges by representing human joints as graph nodes and their connections as edges, enabling simultaneous spatial and temporal dependency modeling. Graph-based Transformer frameworks further incorporate attention for long-range dependencies.
Recent GCN-based approach with SkeleT[10] integrate body, hand, and foot keypoints into a unified graph to capture holistic human dynamics. Although this unified strategy yields strong performance in overall action recognition, it often results in feature dilution: the pooling operations required to handle the scale discrepancy can blur the subtle details of hand motions. This limitation is particularly evident in tasks that demand high precision in hand gesture recognition.
In contrast, our proposed BHaRNet decouples the processing of body and hand features into separate streams. By doing so, we maintain the integrity of fine-grained hand details while still capturing global body dynamics, addressing the key limitation observed in unified approaches (see the experiment section for comparison on hand-intensive actions with the SkeleT model).
@@ II-B Multi-Modal Action Recognition
Multi-modal fusion has emerged as a promising strategy in action recognition, where modalities like RGB, depth, and skeleton data are jointly exploited for robust performance. For example, MMNet [14] is noted for its lightweight design and strong robustness when combining a GCN backbone with an RGB backbone, effectively balancing skeletal and visual information. PoseConv3D[15] uses heatmap-based skeletal learning and bilateral interactions between RGB and skeletal streams, yielding effective multi-modal feature integration.
Figure 3: Shortcut of our Pooling Attention Modules simplifying cross attention mechanism.
@@ II-C Body-Hand Keypoint Integration in Other Fields
Beyond HAR, body–hand integration has been investigated in sign language recognition and keypoint forecasting. In sign language recognition, UNI-SIGN[11] embeds body, hand, and face keypoints into a shared representation of LLM, enabling dynamic modeling of sequential gestures. In parallel, forecasting approaches[12, 13] predict future body–hand coordination, often using specialized alignment or interaction modules to handle the heterogeneity between coarse body dynamics and fine-grained hand movements.
In contrast, our method introduces specialized branches for body and hand, coupled by a streamlined cross-modal attention module—avoiding the need for the shared embedding or the location precision. This design preserves each modality’s unique features and simplifies the fusion process, adapting effectively to fine-grained, context-dependent actions. Moreover, the modular nature of our approach facilitates additional modalities (e.g., RGB) without excessive overhead.
@@ III-A Overview
In this section, we present BHaRNet, a dual-stream network designed to effectively capture both global body dynamics and fine-grained hand articulations. Our model treats the hand as an expert modality, processed independently from the body stream to better capture its unique characteristics.
The overview of our framework is illustrated in Figure LABEL:fig:overview. We first introduce the dual-stream architecture for body and hand, followed by our complementary loss formulation. We then describe a novel cross-attention mechanism that enables controlled feature interaction between the two streams. Finally, we extend our approach to integrate additional RGB data into a multi-modal setting.
Our architecture consists of seperate GCN backbones: one specializing in full-body dynamics and another focused on intricate hand movements. Specifically, a Hand GCN allows for specialized learning of finger movements, whereas a Body GCN focuses on broader pose dynamics. By maintaining seperate initial streams, our model effectively preserves modality-specific features.
@@ III-B Loss functions
A key element of BHaRNet is balancing specialization for each modality with overall synergy. We combine separate classification losses for body and hand with an ensemble loss that considers the fused prediction. Specifically, each modality produces a logit vector($\hat{y_{B}},\hat{y_{H}}$) for its own classification task, leading to individual cross-entropy terms($\mathcal{L}_{B},\mathcal{L}_{H}$). The fused logits, integrated by averaging logit vectors, are treated as a separate classifier with an additional loss term. Our loss implemented in joint training are:
@@
This design ensures that the body stream retains its global pose knowledge while the hand stream focuses on intricate gestures. The ensemble loss also prevents the individual streams from being over-learnt across all action classes, since the gradient stops if one of the streams reason enough on each action class. The logit distributions of the two models are clearly distinctive yielding the simple ensemble that effectively act as a mixture-of-experts (MoE) mechanism.
@@ III-C Cross-Attention for Feature Interaction
Although preserving modality-specific features is essential, interaction between body and hand streams can provide valuable context. We incorporate a cross-attention mechanism enabling feature-level information exchange between modalities. This mechanism manifests through a pooling-attention module and expertized branches.
$\tilde{\mathbf{F}_{B}},\tilde{\mathbf{F}_{H}}=\mathrm{CrossAttn}(\mathbf{F}_{B},\mathbf{F}_{H})\$ (5)
Pooling Attention Module To enchance computational efficiency while retaining rich spatio-temporal information, we introduce a Pooling Attention Module. Instead of applying full cross-attention on high-dimensional feature maps, we perform structured pooling along different axes prior to attention computation.
We decompose feature maps into three different pooled representations, each stream partially pools along two axes (e.g., temporal $T$ and vertical $V$). Body and hand then exchange information via a fast attention-based interaction[20], highlighting informative features from one modality in the other. Afterward, the remaining axis (e.g., instance $I$) is pooled, followed by a fully-connected (FC) layer to produce logits. Applying cross-attention between the partial pooling steps in dimensions $(T,V,I)$, we achieve a balanced trade-off between computational efficiency and rich feature integration.
Let the output of GCN backbone be $\mathbf{F}\in\mathbf{R}^{N\times C\times T\times I\times V}$, where $N$ is the batch size, $C$ the feature channels, $T$ the temporal length, $I$ the number of instances(e.g. body or hand), and $V^{\prime}$ the spatial dimension.
We obtain three pooled features by aggregating different axes(e.g., temporal $T$ and vertical $V$). Let’s denote $\alpha\in\{T,V,I\},\beta,\gamma\in\{T,V,I\}\setminus\{\alpha\}$:
@@
$\mathbf{F}_{B}^{\alpha}=\mathrm{GAP}_{(\alpha)}(\tilde{\mathbf{F}_{B}^{\alpha}}),\mathbf{F}_{H}^{\alpha}=\mathrm{GAP}_{(\alpha)}(\tilde{\mathbf{F}_{H}^{\alpha}})$ (8)
@@ III-D Expertized Branch Model
In this model, the GCN outputs from body and hand each split into two parallel pathways: an Expertized branch that maintains modality-specific features and an Interactive branch that receives cross-attention signals from the complementary modality. The expertized branches pass the raw GCN features directly to GAP and a fully connected layer, thereby preserving each modality’s inherent characteristics. The interactive branches incorporate cross-attention-based fusion before undergoing GAP and classification. We then sum the logits from all four branches ($\mathbf{F}_{1},\mathbf{F}_{2},\mathbf{F}_{3},\mathbf{F}_{4}$) to obtain the final prediction:
TABLE I: Accuracy(%), FLOPs, and Parameter Comparison with State-of-the-Art Methods on NTU RGB+D 60/120. ”-” indicates the experimental results are not provided in the reference, and ”*” for the result based on using public codes
@@
CTR-GCN [8] 92.4 96.8 88.9 90.6 96.5 7.88 5.8
InfoGCN [9] 93.0 97.1 89.8 91.2 97.0 9.96* 9.4
PoseConv3D [15] 94.1 97.1 86.9 90.3 - 31.8 4.0
BlockGCN [6] 93.1 97.0 90.3 91.5 96.9 6.52 5.2
DeGCN (baseline) [21] 93.6 97.4 91.0 92.1 97.2 6.88 5.6
3Mformer[5] 94.8 98.7 92.0 93.8 97.8 58.5 6.7
SkeleT [10] 97.0 99.6 94.6 96.4 97.6 9.6 5.2
@@
Ours (BHaRNet-E, All-ensemble) 96.2 98.8 94.3 95.0 94.6 5.44 5.50
Ours (BHaRNet-E, Expert-only) 96.1 98.7 94.0 94.9 94.6 3.32 2.76
TABLE II: Multi-modal action recognition on NTU60/120, PKU-MMD. ”*” denotes experimental results provided in the other paper.
@@
MMNet (baseline)[14] 96.6 99.1 92.9 94.4 97.4 98.6 89.2 34.2
PoseConv3D (Multi-modal) 97.0 99.6 95.3 96.4 - - 41.8* 31.6*
Ours(BHaRNet-E) 96.2 98.8 94.3 95.0 96.8 97.8 2.72 2.8
Ours(BHaRNet-M) 96.3 99.0 95.1 96.0 96.9 97.9 7.56 16.7
For the Body GCN output $\mathbf{F}_{B}$ and Hand GCN output $\mathbf{F}_{H}$, pipelines of the models are following:
@@
Here, $\mathbf{F}_{1},\mathbf{F}_{2},\mathbf{F}_{3},\mathbf{F}_{4}$ indicate the body-expert, hand-expert, body-interactive, hand-interactive branches, respectively. We also use fast-attention-based cross-attention module as well as Pooling Attention Module. By sharing GCN backbones across branches and applying individual losses only on interactive branches, we prevent the dilution of modality-specific information. Additionally, the complementary loss applied across branches encourages each expertized branch to specialize further, reducing pressure to generalize uniformly across all classes and thereby strengthening each branch’s inherent expertise.
We adopt fast-attention-based cross-attention module to keep computational overhead manageable, as well. In addition, by assigning individual losses only to the interactive branches, we avoid excessive merging of modality-specific features. A complementary loss across branches further encourages each expert branch to focus on what it does best, reducing the need to handle every class uniformly and thereby retaining expert-level cues from body or hand data.
@@ III-E Applications
Light-weighted Inference By selecting only the body- and hand-expertized branches for inference, we derive a simplified model that reduces computational load while retaining the specialized advantages learned from the shared backbone and complementary loss. Though it does not employ cross-attention, its focus on explicit body and hand expertise can offer a pragmatic balance between accuracy and efficiency for scenarios where real-time or low-power constraints are critical.
RGB Integration Inspired by MMNet [14], we incorporate RGB information by treating the body-expert features as guidance for a separate visual stream. This enables our framework to highlight relevant spatio-temporal regions in RGB frames, complementing the skeleton data. In practice, body skeleton data covers global motion, hand skeleton focuses on finer gestures, and RGB frames capture contextual cues such as objects or backgrounds. Together, these data sources enhance each other through cross-attention, resulting in a unified approach suited to a wide range of action recognition tasks.

### v4

@@ I INTRODUCTION
Human Action Recognition (HAR) has garnered significant attention in robotics[wen2023interactive, noisylabels], especially in areas of human-robot interaction and collaboration. The goal of HAR is to automatically identify human actions from video data—a challenging task due to various factors such as complex temporal dependencies, intricate spatial relationships, background noise, and varying camera perspectives. To address these challenges, existing studies have explored diverse data modalities such as RGB video, depth, infrared, audio, and skeleton joints.
Skeleton-based approaches, particularly those using Graph Convolutional Networks (GCNs)[yan2018stgcn, shi2019twoagcn] have shown great promise in efficiently modeling spatial-temporal dynamics by representing skeletal joints as graph nodes and their connections as edges. Recent architectures[wang20233mformer, zhou2024blockgcn, lee2023hierarchically, chen2021ctrgcn, chi2022infogcn] have demonstrated strong performance in action recognition; however, most of these methods primarily focus on full-body motion analysis. This narrow focus limits their ability to capture the fine-grained hand movements essential for differentiating actions such as “Okay sign” or “Victory sign”.
@@
To address these limitations, recent research[yang2024expressive] extends traditional skeleton-based representations by incorporating additional hand and foot keypoints into a unified graph. To manage the increased complexity of this enriched graph, a combined pooling strategy that integrates both node pooling and instance pooling is employed in this approach. While this combined pooling method in an largely-integrated graph significantly enhances efficiency and overall recognition performance, it also results in the loss of subtle hand details, making the hand features vulnerable to blur from larger, dynamic body movements. Among the other related tasks, separating the body and hand into different graph streams is used in sign language recognition[li2025unisign] and action forecasting methods[yan2024forecasting, ding2024expforcasteai] (see Section II for more details).
@@
Each modality is further refined in its dedicated expertized stream. Then, dividing interactive branch and expertized branch enables the two branches to exchange information without using residual connections—thereby preventing conflicts between expertized features and interactive features. Global average pooling is followed by an attention mechanism aggregating body and hand features. Integrating diversely-pooled interactive features, our approach results in a lightweight model with significantly reduced computational complexity.
@@
Inspired by MMNet [bruce2022mmnet], we leverage body skeleton features to guide the learning of RGB features. By treating body features as dynamic weights, our framework enables the RGB stream to focus on complementary contextual information, such as object interactions and environmental cues, enhancing overall recognition performance.
@@ II-A Skeleton-Based Action Recognition
Skeleton-based action recognition has evolved significantly, moving from temporal modeling approaches such as RNNs, to more advanced graph-based methods like GCNs [yan2018stgcn, shi2019twoagcn], CNNs[duan2022poseconv], and Graph-based Transformers [wang20233mformer, duan2023skeletr, do2025skateformer, pang2022igformer, abdelfattah2024maskclr]. RNNs effectively modeled sequential temporal dependencies but struggled with spatial relationships. GCNs addressed these challenges by representing human joints as graph nodes and their connections as edges, enabling simultaneous spatial and temporal dependency modeling. Graph-based Transformer frameworks further incorporate attention for long-range dependencies.
Recent GCN-based approach with SkeleT[yang2024expressive] integrates body, hand, and foot keypoints into a unified graph to capture holistic human dynamics. Although this unified strategy yields strong performance in overall action recognition, it often results in feature dilution: the pooling operations required to handle the scale discrepancy can blur the subtle details of hand motions. This limitation is particularly evident in tasks that demand high precision in hand gesture recognition.
In contrast, our proposed BHaRNet decouples the processing of body and hand features into separate streams. In doing so, we maintain the integrity of fine-grained hand details while still capturing global body dynamics, addressing the key limitation observed in unified approaches (see  Section IV for comparison of hand-intensive actions with the SkeleT model).
Figure 2: Overview of our proposed action recognition model. The model is a dual-stream network consists of Body and Hand expert models, utilizing GCN backbones and cross-attention for enhanced feature fusion.
@@ II-B Multi-Modal Action Recognition
Multi-modal fusion has emerged as a promising strategy in action recognition, where modalities like RGB, depth, and skeleton data are jointly exploited for robust performance. For example, MMNet [bruce2022mmnet] is noted for its lightweight design and strong robustness when combining a GCN backbone with an RGB backbone, effectively balancing skeletal and visual information. PoseConv3D[duan2022poseconv] uses heatmap-based skeletal learning and bilateral interactions between RGB and skeletal streams, yielding effective multi-modal feature integration. A modular nature of our approach via MMNet facilitates additional modalities (e.g., RGB) without excessive overhead.
Figure 3: Shortcut of our Pooling Attention Module simplifying cross attention mechanism.
@@ II-C Body-Hand Keypoint Integration in Other Fields
Beyond HAR, body–hand integration appears in several neighboring domains. For sign‑language recognition, UNI‑SIGN [li2025unisign] embeds body, hand, and face keypoints into a shared LLM‑based representation. Forecasting approaches [yan2024forecasting, ding2024expforcasteai] predict future body–hand coordination via specialized alignment modules, and REWIND [lee2025rewind] demonstrates the value of accurate body–hand keypoints for egocentric pose estimation. Distinct from these methods, we preserve body and hand as separate experts and employ lightweight cross‑attention, avoiding the need for the shared embedding or the location precision.
II-D Fine‑grained Hand Interaction
Recent studies emphasize that preserving subtle finger articulations is crucial for understanding or synthesizing hand‑centric behavior. On the recognition side [shamil2024utility, kwon2021h2o, cho2023transformer], accurate 3D or dual‑hand signals improved fine‑grained classification. Other studies [li2025latenthoi, woo2024hand] model detailed hand–object contacts, while two‑hand reconstruction approaches [lee2023im2hands, kim2024bitt] recover realistic geometry or texture from monocular input. These lines of research underline the importance of preserving fine‑grained hand cues.
@@ III-A Overview
In this section, we present BHaRNet, a dual-stream network designed to effectively capture both global body dynamics and fine-grained hand articulations. Our model treats the hand as an expert modality, processed independently ofhe body stream to better capture its unique characteristics. The overview of our framework is illustrated in Figure 2.
We first introduce the dual-stream architecture for body and hand, followed by our complementary loss formulation. We then describe a novel cross-attention mechanism that enables controlled feature interaction between the two streams. Finally, we extend our approach to integrate additional RGB data into a multi-modal setting.
Our architecture consists of separate GCN backbones: one specializing in full-body dynamics and another focused on intricate hand movements. Specifically, a hand GCN allows for specialized learning of finger movements, whereas a body GCN focuses on broader pose dynamics. By maintaining separate initial streams, our model effectively preserves modality-specific features.
@@ III-B Loss functions
A key element of BHaRNet is balancing the specialization of each modality with the overall synergy. We combine separate classification losses for body and hand with complementary loss that considers the fused prediction. Specifically, each modality produces a logit vector($\hat{y_{B}},\hat{y_{H}}$) for its own classification task, leading to individual cross-entropy terms($\mathcal{L}_{B},\mathcal{L}_{H}$) with the ground-truth label($l$). The fused logits, integrated by averaging logit vectors, are treated as a separate classifier with an additional loss term. Our loss implemented in joint training are:
@@
This design ensures the body stream retains its global pose knowledge, while the hand stream focuses on intricate gestures. The complementary loss also prevents the individual streams from being over-learned across all action classes, since the gradient stops if one of the streams reason enough on each action class. The logit distributions of the two models are clearly distinctive yielding the simple ensemble that effectively acts as a mixture-of-experts (MoE) mechanism.
@@ III-C Cross-Attention for Feature Interaction
While preserving modality-specific features is essential, interaction between body and hand streams can provide valuable context. We incorporate a cross-attention mechanism enabling feature-level information exchange between modalities. This mechanism manifests through a pooling attention module.
$\tilde{\mathbf{F}_{B}},\tilde{\mathbf{F}_{H}}=\mathrm{CrossAttn}(\mathbf{F}_{B},\mathbf{F}_{H})\\$ (5)
Pooling Attention Module. To enchance computational efficiency while retaining rich spatio-temporal information, we introduce a Pooling Attention Module. Instead of applying full cross-attention on high-dimensional feature maps, we perform structured pooling along different axes prior to attention computation.
We decompose feature maps into three different pooled representations, each stream partially pools along two axes (e.g. temporal $T$ and vertical $V$). Body and hand then exchange information via a fast attention-based interaction[choromanski2020performeratt], highlighting informative features from one modality in the other. Afterward, the remaining axis (e.g. instance $I$) is pooled, followed by a fully-connected (FC) layer to produce logits. Applying cross-attention between the partial pooling steps in dimensions $(T,V,I)$, we achieve a balanced trade-off between computational efficiency and rich feature integration.
Let the output of GCN backbone be $\mathbf{F}\in\mathbf{R}^{N\times C\times T\times I\times V}$, where $N$ is the batch size, $C$ the feature channels, $T$ the temporal length, $I$ the number of instances(body or hand), and $V$ the spatial dimension. For $\alpha\in\{T,V,I\},\beta,\gamma\in\{T,V,I\}\setminus\{\alpha\}$:
@@
[…]
@@ III-D Expertized Branch Model
In this model, the GCN outputs from body and hand each split into two parallel pathways: an Expertized branch that maintains modality-specific features and an Interactive branch that receives cross-attention signals from the complementary modality. The expertized branches($\mathbf{F}_{1},\mathbf{F}_{2}$) pass the raw GCN features directly to GAP and FC layer, thereby preserving each modality’s inherent characteristics. The interactive branches($\mathbf{F}_{3},\mathbf{F}_{4}$) incorporate cross-attention-based fusion before undergoing GAP and classification. We then sum the logits from all four branches to obtain the final prediction:
TABLE I: Accuracy(%), FLOPs, and Parameter Comparison with State-of-the-Art Methods on NTU RGB+D 60/120. “-” indicates the experimental results are not provided in the reference, and “*” for the result based on using public codes. Best results are highlighted as first and second.
@@
CTR-GCN [chen2021ctrgcn] 92.4 96.8 88.9 90.6 96.5 7.9 5.8
InfoGCN [chi2022infogcn] 93.0 97.1 89.8 91.2 97.0 10.0* 9.4
PoseConv3D [duan2022poseconv] 94.1 97.1 86.9 90.3 - 31.8 4.0
BlockGCN [zhou2024blockgcn] 93.1 97.0 90.3 91.5 96.9 6.5 5.2
DeGCN (baseline) [myung2024degcn] 93.6 97.4 91.0 92.1 97.2 6.9 5.6
3Mformer[wang20233mformer] 94.8 98.7 92.0 93.8 97.8 58.5 6.7
SkeleT [yang2024expressive] 97.0 99.6 94.6 96.4 97.6 9.6 5.2
Ours (BHaRNet-E, Expert-only) 96.1 98.7 94.0 94.9 94.6 3.3 2.8
Ours (BHaRNet-E, All-ensemble) 96.2 98.8 94.3 95.0 94.6 5.44 5.5
@@
TABLE II: Multi-modal action recognition on NTU60/120, PKU-MMD. “*” denotes experimental results provided in the other paper.
@@
MMNet (baseline)[bruce2022mmnet] 96.6 99.1 92.9 94.4 97.4 98.6 89.2 34.2
PoseConv3D[duan2022poseconv] 97.0 99.6 95.3 96.4 - - 41.8* 31.6*
Ours(BHaRNet-M) 96.3 99.0 95.1 96.0 96.9 97.9 7.6 16.7
@@
Here, $\mathbf{F}_{1},\mathbf{F}_{2},\mathbf{F}_{3},\mathbf{F}_{4}$ indicates the branch of expert body, expert hand, interactive body, interactive hand, respectively. We also use a fast-attention-based cross-attention module to keep computational overhead manageable. To explicitly encourage expertized branches to specialize further, we modify our individual loss formulation:
$\mathcal{L}_{B}=\mathrm{CE}\bigl(\hat{y_{3}},\,l\bigr),\quad\mathcal{L}_{H}=\mathrm{CE}\bigl(\hat{y_{4}},\,l\bigr)$ (13)
In addition, a complementary loss is applied across all branches, as defined in Section III-B. By assigning individual losses only to the interactive branches, we avoid excessive merging of modality-specific features. The complementary loss further encourages each expert branch to focus on its inherent strengths, reducing the need to handle every class uniformly and thereby retaining expert-level cues from body or hand data.
@@ III-E Applications
Light-weighted Inference. By selecting only the body- and hand-expertized branches for inference, we derive a simplified model that reduces computational load while retaining the specialized advantages learned from the shared backbone and complementary loss. Though it does not employ cross-attention, its focus on explicit body and hand experts can offer a pragmatic balance between accuracy and efficiency for scenarios where low-power constraints are critical.
RGB Integration. Inspired by MMNet [bruce2022mmnet], we incorporate RGB information by treating body-expert features as guidance for a separate video stream. This enables our framework to highlight relevant spatio-temporal regions in RGB frames, complementing the skeleton data. In practice, body skeleton data covers global motion, hand skeleton focuses on finer gestures, and RGB frames capture contextual cues such as objects or backgrounds. Together, these data sources enhance each other through cross-attention, resulting in a unified approach suited to a wide range of action recognition tasks.